# NNN: Next-Generation Neural Networks for Marketing Mix Modeling

### v2

@@ 12 Experiments
When computing sale-prediction error rates, we always roll-up predictions to the national-level, like in Meridian MMM [35]. We evaluate the MAPE and R2 on three datasets: train, validation, and test. The train dataset is used to fit the parameters, while the other two datasets contain hold-out data. In all experiments, the test dataset represents data outside the temporal training window (i.e., from dates greater than the last training date), while the validation data is sampled from within the same training window (although this data is not use during model fitting). Both datasets contain unseen data, and model accuracy measured on either is a proxy for how well the model generalizes, however, generally the test set generalizes with more difficulty, because its data represent samples from a different time period than the data used to fit the model.
@@ 12.1 Simulated Data
We compare our method, NNN, against a traditional MMM on both forms of the synthetic dataset. We measure the predictive error rates on validation and test sets. Using the models fit only on the train dataset, we calculate the sales attribution of each media channel. Meridian is fit to the data using the default parameters and priors. NNN is also fit using its default parameters. Due to the software setup in Meridian, we trained two models: the first model is trained using the train data and evaluated on the validation data, while the second model is trained on the train and validation data and evaluated on the test data. Due to this, we show a pair of values that corresponding to training metrics for MMM. We compute the mixfor each channel and each model. For NNN, we train two versions of the model: one using only the Sales as the target and another using both Search and sales as the target. Both models are trained without channel mixing for 5k steps. For the Search and sales model, we train for an additional 5k steps using balancing coefficient of .99, and we report the mixusing both the traditional method and the autoregressive (AR) unrolling method.
@@
Tables 1 and 2 show the predictive errors for models trained on the Low Variance and High Variances datasets, respectively. For both datasets, NNN outperforms MMM on both holdout sets. For the Low Variance dataset, both models achieve good MAPE metrics on all data splits, however NNN does a much better job at explaining the variance of the data, as measured by R2. Interestingly, NNN achieves MAPE values that may be approaching values generated by the synthetic noise, without overfitting to the train dataset. For the High Variance dataset, NNN greatly outperforms MMM in both MAPE and R2. Most of the variance in the sales can be measured using a subset of the Search terms (i.e., the high-impact terms), which NNN appears to correctly parse using the embeddings, while MMM, which only uses the Search volume, is unable to differentiate between the high-impact and low-impact terms.
@@
For the both datasets, Meridian provides decent estimates across all channels, however NNN outperforms MMM for the Search and YouTube channels. Both MMM and NNN under predict the Search Ads mix on both datasets, however MMM outperformed NNN. Both MMM and NNN are closer to the Direct values than the Total values for the Search and YouTube channels, as expected. When utilizing AR unrolling, we see that NNN properly moves sales from Search to YouTube, however it slightly underestimates the YouTube contribution (39.65%) for the High Variance Dataset, and it slightly overestimates for the Low Variance Dataset (55.10%). Interestingly, it also adjusts the Search Ads mix estimates. We leave further investigating to future work [^8] Particularly in the area of more structural Search heads and a robust method for determining the balancing coefficient., although we are excited about the promise of this direction.
@@
This finding suggests that minimizing predictive error on appropriately chosen OOD datasets is a promising strategy for identifying NNN models that also yield accurate channel attributions. This perspective contrasts somewhat with discussions in traditional MMM literature [7], which often caution against relying solely on statistical fit metrics for model selection due to potential issues like model multiplicity on limited or non-representative data. We hypothesize that NNN’s success with this OOD-driven approach stems from its evaluation across multiple hold-out sets, particularly the use of a test dataset drawn from a temporally distinct period from the training data. This provides a more challenging and informative signal for generalization compared to evaluations restricted to in-sample data or temporally similar validation sets. Furthermore, the sparsity induced by L1 regularization appears particularly beneficial for managing the large parameter spaces inherent in neural network architectures, especially within the often data-constrained context of MMM.
From a practical standpoint, this relationship between OOD performance and attribution accuracy potentially streamlines the model selection workflow compared to traditional MMM practices. Instead of potentially lengthy, iterative cycles of selecting priors, retraining, and evaluating based on parameter plausibility and fit, our approach centers on tuning a single key hyperparameter ($\lambda$) via a readily parallelizable grid search. Models can be primarily selected based on achieving optimal predictive performance on OOD validation/test sets. While this data-driven selection forms the core, we recommend that chosen candidate models still undergo verification through attribution analysis and the simulation-based probing discussed previously and below, ensuring the selected model behaves reasonably beyond aggregate predictive metrics.
12.1.1 Model Application: Simulation and Analysis
@@
Marketing Pause We simulate the effect of a YouTube pause by defining a twenty-week period to set the YouTube marketing volume to zero. We then compute model inference two ways: using autoregressive unrolling and using the standard inference with the observed values of Search. This simulation is shown in Figure 10. The estimated sales for both pause experiments are smaller than the estimated sales using the observed marketing and search data (i.e., no pause), however the effect of the pause is much more pronounces when accounting for the effect YouTube had on Search, and the pause effect lasts much longer.
@@
This analysis framework can be applied to any channel for which specific embeddings (creatives, keywords, etc.) are available. It provides valuable insights into creative or keyword effectiveness as an intrinsic artifact of the MMM training process, complementing traditional measurement methods and serving as an additional tool for model validation. Furthermore, it exemplifies the broader capability of using the trained NNN model for diverse simulation-based analyses beyond these specific examples.
@@ 12.2 Real World Marketing Data
First, we compare Bayesian MMM against our method on a real world dataset. Both models were trained at a regional level (4 regions were used, as defined by U.S. Census), and the data was at a weekly granularity.
Previously, an MMM was fit for our end user, and we use the formerly informed priors to fit this version of the model. We used standard fitting methods used in practice. The time-variant effect was applied to this MMM (at the end of each month we insert one coefficient). In order to give a fair comparison on the test set, which appears after all training and validation points, we allowed the MMM to fit two additional points to the spline during this period. [^9] This effectively allowed the MMM to observe the holdout set during fitting, but it was the only way we could see giving a somewhat comparable estimation on data for test the timeframe.. We used 7 chains with 1000 burn-ins, and all the channels converged during optimization. We only tuned the standard deviations of the priors while looking at model diagnostics and convergence (i.e., if the baseline had a negative contribution, we retrained). Similar to the simulated datasets, we followed a two-model training procedure for validation and test evaluations.
@@
We notice that the traditional MMM does a good job at predicting out of sample data during the training period (validation MAPE 20% and R2 .77), however it suffers greatly during the test period (189% MAPE, -90 R2). The NNN Baseline generalizes well in both the validation data (5.3% MAPE, .99 R2) and the test data (16.2% MAPE, .92 R2), achieving better errors than MMM. When the additional channels are included, the validation and test errors are further reduced, demonstrating that NNN can make effective use of the additional channels using the same format of data traditionally available in MMM.
@@ 12.2.1 Additional Experiments
13 Discussion: NNN in the Context of Traditional MMM
The introduction of NNN presents a significant departure from traditional MMM methodologies. By leveraging the representational power of neural networks, Transformer architectures, and rich embeddings of marketing channels alongside organic signals like Google Search, NNN offers a novel framework for modeling marketing effectiveness. This section compares NNN to traditional MMM approaches across several key dimensions, highlighting differences in methodology and capability.
Modeling Foundations. Traditional MMMs typically employ established statistical models, often based on linear regression, incorporating transformations like adstock and Hill functions to capture non-linearities and carryover effects. NNN, in contrast, utilizes the inherent non-linear function approximation capabilities of deep neural networks and the sequence modeling power of Transformers and self-attention. This allows NNN to potentially capture more complex interactions and temporal dependencies directly from the data.
Furthermore, MMMs primarily rely on scalar time-series representations of marketing inputs (e.g., weekly spend or impressions). NNN moves beyond this using the advances in representation learning, by ingesting rich, high-dimensional embeddings. These embeddings integrate both quantitative metrics (like impressions, encoded in the embedding’s magnitude) and qualitative aspects (like ad creative content or search query semantics, encoded in the embedding’s direction). With this semantic information, the hope is that NNN can distinguish between, for example, a million impressions of a brilliant ad versus a million impressions of a dud. The stark performance difference observed on the High Variance simulated dataset, where the semantic nuance was deliberately included, underscores the value of this approach.
@@ 13 Discussion: NNN in the Context of Traditional MMM
NNN, on the other hand, largely eschews this kind of parameter-level prior specification. While its network weights are initialized before training, the methods typically used—standard schemes like Glorot and He initialization[14, 20]—are chosen primarily to ensure stable gradient flow and effective optimization, rather than to encode any specific belief about the final value of a given weight parameter [^10] Although there are cases where special care is taken, such as the identity matrix initialization for the channels matrix in our attention mechanism.. In general, there are too many parameters, interacting in too complex a way, for individual weight priors to be easily interpretable or specifiable in the BMMM sense. Instead, in NNN, prior beliefs are encoded more implicitly, baked directly into the architecture and computational structure of the model itself. For example, when designing our Transformer’s attention mechanism, the use of a causal mask rigidly enforces the prior belief that future events cannot influence past predictions. Similarly, the deliberate removal of layer normalization components within the Transformer stems from the prior understanding that the magnitude of our embedding vectors carries crucial information about marketing volume, information that normalization might otherwise discard. These architectural choices act as strong structural priors, shaping the function space the model explores, even if individual parameters aren’t given explicit Bayesian priors.
@@
Beyond these architectural choices, the optimization process itself introduces another form of implicit prior via regularization. Our NNN framework employs an L1 penalty added to the loss function during training. This L1 term actively encourages sparsity by pushing many model parameters towards zero. This effectively embeds a belief in sparsity within the learning process itself – an assumption that, within the potentially vast parameter space of the neural network, many connections or features are ultimately unnecessary for explaining the data. However, this sparsity-inducing prior is typically chosen more for its beneficial regularization effects—combating overfitting in high dimensions and aiding model selection via OOD performance-than for encoding specific, quantifiable beliefs about the likely value of any individual parameter, unlike the explicit distributional priors of BMMM.
@@
Model Uncertainty. Bayesian MMMs offer a significant advantage in inherently quantifying uncertainty through posterior distributions for parameters and predictions. By using priors, and MCMC sampling, they provide not just point estimates, but posterior distributions for model parameters and predictions. This furnishes MMMs with a statistically grounded measure of confidence. NNN is trained by providing point estimates of sales and media channel values, then using gradient-based optimization moves the point estimate of the parameters. This provides neither uncertainty estimates of the model parameters nor the predictions. Because the NNN model is a neural network, providing distributions over the parameters won’t make much practical difference over the current point estimates: these parameters are usually not inspected. However, providing uncertainty distributions for the predictions is useful, and we look to techniques like Monte Carlo Dropout as a viable path for incorporating uncertainty estimation in future work.
@@
Capturing Temporal Dynamics: Seasonality, Baselines, and Long-Term Effects. Handling temporal patterns is critical in marketing modeling. Traditional MMMs often address seasonality using explicit indicator variables or spline knots, while carryover is modeled with parametric adstock functions, which typically struggle to capture effects beyond a few months. Baseline demand is often represented by simple trend components.
@@
NNN takes a different approach by leveraging rich, dynamic signals and its sequence modeling architecture. We rely on organic signals, particularly Google Search embeddings, to implicitly capture complex baseline fluctuations, including seasonality. As prior work suggests, Search behavior naturally reflects weekly, seasonal, and holiday-related patterns. NNN’s neural networks can parse this relevant information directly from the Search embeddings, reducing the need for extensive seasonal feature engineering. Moreover, by explicitly modeling the influence of marketing on intermediary signals like Search and employing a Transformer architecture with its inherent attention mechanism over long sequences, NNN has the potential to capture marketing effects unfolding over much longer time horizons than traditional adstock allows. Our simulation experiments, particularly the YouTube pause analysis (Figure 10), provide evidence for this capability.
@@
Attribution and Interpretation. Both NNN and MMMs generally resort to counterfactual simulations—running the model with and without a specific channel’s activity—to estimate its contribution. However, traditional MMMs, often provide direct interpretability through parameter estimates (e.g., coefficients, decay rates, saturation parameters) within well-defined functional forms.
@@
NNN, like many deep learning models, faces challenges in direct parameter interpretability. However, we believe its strength lies in modeling the dynamic interplay between channels and its ability to perform analyses like the creative/keyword analysis. By probing the model’s response to specific, meaningful inputs (like different ad embeddings), we can gain valuable insights into the learned relationships and drivers of effectiveness, offering a different but complementary form of interpretability. Furthermore, modeling intermediary channels like Search allows for potentially more nuanced attribution, distinguishing between direct effects on sales and indirect effects mediated through organic demand shifts.
@@
NNN, a machine learning-forward approach, leans more heavily on objective, out-of-distribution (OOD) predictive performance as its North Star. Metrics like MAPE and R2 calculated on held-out validation set and, crucially, a temporally distinct test set are the primary criteria for selecting hyperparameters, particularly the L1 regularization strength which governs model sparsity. The rationale, supported by our experiments, is that models generalizing well to unseen, OOD data are more likely to have captured robust underlying patterns and thus yield more reliable attribution estimates, even if their internal parameters are complex or non-intuitive. This allows exploration of highly parameterized models that might be discarded in traditional MMM workflows. This data-driven selection process, which can involve parallel model fitting, can be significantly faster than iterative prior adjustment. However, given the novelty of the approach and the inherent risks of complex models, we strongly recommend supplementing OOD performance checks with the kinds of simulation-based probing (attribution, creative analysis) discussed earlier as a crucial vetting step.
Summary. NNN introduces a different set of trade-offs compared to traditional MMMs. It seems to excel in leveraging rich embedding representations, modeling complex non-linearities and long-term temporal dynamics via its neural and Transformer components, and utilizing predictive performance on holdout data for model selection. Traditional approaches, particularly Bayesian MMMs, currently hold advantages in direct parameter interpretability and built-in uncertainty quantification. NNN represents a promising direction, especially suited for scenarios rich in contextual data where capturing complex interactions and long-range effects is paramount, though further validation and development.

### v3

@@ 12 Experiments
When computing sales-prediction error rates, we always roll-up predictions to the national-level to make results easier to interpret. We evaluate the MAPE and R2 on three datasets: train, validation, and test. The train dataset is used to fit the parameters, while the other two datasets contain hold-out data. In all experiments, the test dataset represents data outside the temporal training window (i.e., from dates greater than the last training date), while the validation data is sampled from within the same training window (although this data is not use during model fitting). Both datasets contain unseen data, and model accuracy measured on either is a proxy for how well the model generalizes, however, generally the test set generalizes with more difficulty, because its data represent samples from a different time period than the data used to fit the model.
@@ 12.1 Simulated Data
We compare our method, NNN, against a our MMM baseline on both forms of the synthetic dataset. We measure the predictive error rates on validation and test sets. Using the models fit only on the train dataset, we calculate the sales attribution of each media channel. The baseline MMM is fit to the data using the default parameters and priors. NNN is also fit using its default parameters. We trained two models: the first model is trained using the train data and evaluated on the validation data, while the second model is trained on the train and validation data and evaluated on the test data. Thus we show a pair of values that corresponding to training metrics for MMM. We compute the mixfor each channel and each model. For NNN, we train two versions of the model: one using only the Sales as the target and another using both Search and sales as the target. Both models are trained without channel mixing for 5k steps. For the Search and sales model, we train for an additional 5k steps using balancing coefficient of .99, and we report the mixusing both the traditional method and the autoregressive (AR) unrolling method.
@@
Tables 1 and 2 show the predictive errors for models trained on the Low Variance and High Variances datasets, respectively. It should be noted that MMMs are not usually evaluated primarily in terms of their predictive error since they are not predictive models; instead MMMs are explanatory models and/or designed for causal inference. In the NNN framework that we propose in this paper, however, the NNNs can be treated as predictive models so we adopt that comparison here.
For both datasets, NNN outperforms our baseline MMM on both holdout sets in terms of R2 and MAPE. For the Low Variance dataset, both models achieve good MAPE metrics on all data splits, however NNN does a much better job at explaining the variance of the data, as measured by R2. Interestingly, NNN achieves MAPE values that may be approaching values generated by the synthetic noise, without overfitting to the train dataset. For the High Variance dataset, NNN outperforms MMM in both MAPE and R2 as well. Most of the variance in the sales can be measured using a subset of the Search terms (i.e., the high-impact terms), which NNN appears to correctly parse using the embeddings, while MMM, which only uses the Search volume, is unable to differentiate between the high-impact and low-impact terms.
@@
For both datasets, the baseline MMM provides decent estimates across all channels. The NNN is closer to the ground truth for YouTube and for Search, while the MMM is closer to ground truth for Search Ads. Both models under-predict the attribution of Search Ads, and both models are closer to the Direct values than the Total values for the Search and YouTube channels, as expected. When utilizing AR unrolling, we see that NNN properly moves sales from Search to YouTube, however it slightly underestimates the YouTube contribution (39.65%) for the High Variance Dataset, and it slightly overestimates for the Low Variance Dataset (55.10%). Interestingly, it also adjusts the Search Ads mix estimates. We leave further investigating to future work [^8] Particularly in the area of more structural Search heads and a robust method for determining the balancing coefficient., although we are excited about the promise of this direction.
@@
This finding suggests that minimizing predictive error on appropriately chosen OOD datasets is a promising strategy for identifying NNN models that also yield accurate channel attributions. This perspective contrasts somewhat with the way that MMMs are generally built [7], because MMMs are generally explanatory or causal models rather than predictive models. Thus, within that framework, MMM authors often caution against relying solely on statistical fit metrics for model selection. We hypothesize that NNN’s success with this OOD-driven approach stems from its evaluation across multiple hold-out sets, particularly the use of a test dataset drawn from a temporally distinct period from the training data. This provides a more challenging and informative signal for generalization compared to evaluations restricted to in-sample data or temporally similar validation sets. Furthermore, the sparsity induced by L1 regularization appears particularly beneficial for managing the large parameter spaces inherent in neural network architectures, especially within the often data-constrained context of MMM.
From a practical standpoint, this relationship between OOD performance and attribution accuracy potentially streamlines the model selection workflow. Instead of potentially lengthy, iterative cycles of selecting priors, retraining, and evaluating based on parameter plausibility and fit, our approach centers on tuning a single key hyperparameter ($\lambda$) via a readily parallelizable grid search. Models can be primarily selected based on achieving optimal predictive performance on OOD validation/test sets. While this data-driven selection forms the core, we recommend that chosen candidate models still undergo verification through attribution analysis and the simulation-based probing discussed previously and below, ensuring the selected model behaves reasonably beyond aggregate predictive metrics. As we will discuss later in this paper, future validation and research will be required in order to assess the robustness and quantify the uncertainty distribution of measurements made under the NNN framework.
12.1.1 Model Application beyond the MMM framework: Simulation and Analysis
@@
Predictions As [28] demonstrated, the representation of the weighted search embedding and the search volume in these models can be incredibly effective prediction tools. We expect the same results would hold here, with the addition of including the effects of media on sales in addition to organic search.
Marketing Pause We simulate the effect of a YouTube pause by defining a twenty-week period to set the YouTube marketing volume to zero. We then compute model inference two ways: using autoregressive unrolling and using the standard inference with the observed values of Search. This simulation is shown in Figure 10. The estimated sales for both pause experiments are smaller than the estimated sales using the observed marketing and search data (i.e., no pause), however the effect of the pause is much more pronounced when accounting for the effect YouTube had on Search, and the pause effect lasts much longer.
@@
This analysis framework can be applied to any channel for which specific embeddings (creatives, keywords, etc.) are available. When using such channels, it provides valuable insights into creative or keyword effectiveness in parallel to media measurement for these channels. However, NNN does require these embeddings in order to generate these additional insights and use the content or the context of the ads for its modeling (and for many traditional media channels like TV these will not be readily available). We have not yet done comprehensive research on the effectiveness of this framework with purely scalar channels, especially at the scale of dozens of channels which is common for MMMs.
@@ 12.2 Real World Marketing Data
First, we compare our baseline MMM against our NNN method on a real world dataset. Both models were trained at a regional level (4 regions were used, as defined by U.S. Census), and the data was at a weekly granularity.
Previously, an MMM was fit for our end user, and we use the formerly informed priors to fit this version of the model. We used standard fitting methods used in practice. A time-variant effect was applied to this MMM (at the end of each month we insert one coefficient). In order to give a fair comparison on the test set, which appears after all training and validation points, we allowed the MMM to fit two additional points to the baseline during this period. [^9] This effectively allowed the MMM to observe the holdout set during fitting, but it was the only way we could see giving a somewhat comparable estimation on data for test the timeframe.. We used 7 chains with 1000 burn-ins, and all the channels converged during optimization. We only tuned the standard deviations of the priors while looking at model diagnostics and convergence (i.e., if the baseline had a negative contribution, we retrained). Similar to the simulated datasets, we followed a two-model training procedure for validation and test evaluations.
@@
We notice that the baseline MMM does a good job at predicting out of sample data during the training period (validation MAPE 20% and R2 .77), however it suffers during the test period (189% MAPE, .90 R2). The NNN Baseline generalizes well in both the validation data (5.3% MAPE, .99 R2) and the test data (16.2% MAPE, .92 R2), achieving better errors than MMM. When the additional channels are included, the validation and test errors are further reduced, demonstrating that NNN can make effective use of the additional channels using the same format of data traditionally available in MMM.
@@ 12.2.1 Additional Experiments
13 Discussion: NNN in the Context of MMM
The introduction of NNN presents a significant departure from general MMM methodologies. By leveraging the representational power of neural networks, Transformer architectures, and rich embeddings of marketing channels alongside organic signals like Google Search, NNN offers a novel framework for modeling marketing effectiveness. This section compares NNN to standard MMM approaches across several key dimensions, highlighting differences in methodology and capability.
Modeling Foundations. NNN sits outside the established statistical frameworks used for MMMs. As we mentioned above, while some of its functionality overlaps with the types of media measurement that MMMs provide, it does not necessarily replace MMM in real-world contexts. As we discuss below, NNNs require significant additional research and validation.
MMMs typically employ established statistical models, often based on linear regression, incorporating transformations like adstock and Hill functions to capture non-linearities and carryover effects. NNN, in contrast, NNN utilizes the inherent non-linear function approximation capabilities of deep neural networks and the sequence modeling power of Transformers and self-attention. This allows NNN to potentially capture more complex interactions and temporal dependencies directly from the data, although again more research is required here.
Furthermore, MMMs primarily rely on scalar time-series representations of marketing inputs (e.g., weekly spend or impressions). NNN uses these data as well, but supplements them with rich, high-dimensional embeddings. These embeddings integrate both quantitative metrics (like impressions, encoded in the embedding’s magnitude) and qualitative aspects (like ad creative content or search query semantics, encoded in the embedding’s direction). With this semantic information, the hope is that NNN can distinguish between, for example, a million impressions of a brilliant ad versus a million impressions of a dud. The performance difference observed on the High Variance simulated dataset, where the semantic nuance was deliberately included, underscores the value of this approach.
@@ 13 Discussion: NNN in the Context of MMM
NNN, on the other hand, largely eschews this kind of parameter-level prior specification. While its network weights are initialized before training, the methods typically used—standard schemes like Glorot and He initialization[14, 20]—are chosen primarily to ensure stable gradient flow and effective optimization, rather than to encode any specific belief about the final value of a given weight parameter [^10] Although there are cases where special care is taken, such as the identity matrix initialization for the channels matrix in our attention mechanism.. In general, there are too many parameters, interacting in too complex a way, for individual NNN weights to be easily interpretable or specifiable in a Bayesian sense. Instead, in NNN, prior beliefs are encoded more implicitly, baked directly into the architecture and computational structure of the model itself. For example, when designing our Transformer’s attention mechanism, the use of a causal mask rigidly enforces the prior belief that future events cannot influence past predictions. Similarly, the deliberate removal of layer normalization components within the Transformer stems from the prior understanding that the magnitude of our embedding vectors carries crucial information about marketing volume, information that normalization might otherwise discard. These architectural choices act as strong structural priors, shaping the function space the model explores, even if individual parameters aren’t given explicit Bayesian priors.
@@
Beyond these architectural choices, the optimization process itself introduces another form of implicit prior via regularization. Our NNN framework employs an L1 penalty added to the loss function during training. This L1 term actively encourages sparsity by pushing many model parameters towards zero. This effectively embeds a belief in sparsity within the learning process itself – an assumption that, within the potentially vast parameter space of the neural network, many connections or features are ultimately unnecessary for explaining the data. However, this sparsity-inducing prior is typically chosen more for its beneficial regularization effects—combating overfitting in high dimensions and aiding model selection via OOD performance-than for encoding specific, quantifiable beliefs about the likely value of any individual parameter, unlike the explicit distributional priors of Bayesian models.
@@
Model Uncertainty. Bayesian MMMs offer a significant advantage in inherently quantifying uncertainty through posterior distributions for parameters and predictions. By using priors, and MCMC sampling, they provide not just point estimates, but posterior distributions for model parameters and predictions. This furnishes Bayesian MMMs with a statistically grounded measure of confidence. In constrast NNN is trained by providing point estimates of sales and media channel values, then using gradient-based optimization moves the point estimate of the parameters. This provides neither uncertainty estimates of the model parameters nor the predictions. Because the NNN model is a neural network, providing distributions over the parameters won’t make much practical difference over the current point estimates: these parameters are usually not inspected. However, providing uncertainty distributions for the predictions is useful, and we look to techniques like Monte Carlo Dropout as a viable path for incorporating uncertainty estimation in future work.
@@
Capturing Temporal Dynamics: Seasonality, Baselines, and Long-Term Effects. Handling temporal patterns is critical in marketing modeling. Many MMMs address seasonality using explicit indicator variables or spline knots, while carryover is modeled with parametric adstock functions, which typically struggle to capture effects beyond a few months. Baseline demand is often represented by simple trend components.
@@
NNN takes a different approach by leveraging rich, dynamic signals and its sequence modeling architecture. We rely on organic signals, particularly Google Search embeddings, to implicitly capture complex baseline fluctuations, including seasonality. As prior work suggests, Search behavior naturally reflects weekly, seasonal, and holiday-related patterns. NNN’s neural networks can parse this relevant information directly from the Search embeddings, reducing the need for extensive seasonal feature engineering. Moreover, by explicitly modeling the influence of marketing on intermediary signals like Search and employing a Transformer architecture with its inherent attention mechanism over long sequences, NNN has the potential to capture marketing effects unfolding over much longer time horizons than traditional adstock allows. Our simulation experiments, particularly the YouTube pause analysis (Figure 10), provide evidence for this capability, although further research and validation is needed.
@@
Attribution and Interpretation. Both NNN and MMMs generally resort to counterfactual simulations—running the model with and without a specific channel’s activity—to estimate its contribution. However MMMs also can provide direct interpretability through parameter estimates (e.g., coefficients, decay rates, saturation parameters) within well-defined functional forms.
@@
NNN, like many deep learning models, faces challenges in direct parameter interpretability. However, we believe its strength lies in modeling the dynamic interplay between selected channels and its ability to perform analyses like the creative/keyword analysis. By probing the model’s response to specific, meaningful inputs (like different ad embeddings), we can gain valuable insights into the learned relationships and drivers of effectiveness, offering a different but complementary form of interpretability. Furthermore, modeling intermediary channels like Search allows for potentially more nuanced attribution, distinguishing between direct effects on sales and indirect effects mediated through organic demand shifts.
@@
NNN, a machine learning-forward approach, leans more heavily on objective, out-of-distribution (OOD) predictive performance as its North Star. Metrics like MAPE and R2 calculated on held-out validation set and, crucially, a temporally distinct test set are the primary criteria for selecting hyperparameters, particularly the L1 regularization strength which governs model sparsity. The rationale, supported by our experiments, is that models generalizing well to unseen, OOD data are more likely to have captured robust underlying patterns and thus yield more reliable attribution estimates, even if their internal parameters are complex or non-intuitive. This data-driven selection process, which can involve parallel model fitting, can be significantly faster than iterative prior adjustment.
However, given the novelty of the approach, the limited set of examples examined so far in this paper, and the inherent risks of complex models, we strongly recommend supplementing these OOD performance checks with business judgment and context. This can be introduced through the kinds of simulation-based probing (attribution, creative analysis) discussed earlier as a crucial vetting step.
Summary. As it covers a slightly different set of potential use cases, NNN introduces a different set of trade-offs compared to MMM. It seems to excel in leveraging rich embedding representations, modeling complex non-linearities and long-term temporal dynamics via its neural and Transformer components, and utilizing predictive performance on holdout data for model selection. Current MMM approaches, particularly Bayesian MMMs, hold advantages in direct parameter interpretability and built-in uncertainty quantification. NNN represents a promising research direction, especially suited for scenarios rich in contextual data where capturing complex interactions and long-range effects is paramount, though further validation and development is needed.